OpenStudy (anonymous):

Write the equation of the line perpendicular to 9x+2y=18 and passes through the point (1,4)

OpenStudy (anonymous):

This is what I got... 9x+2y=18 2y = -9x + 18 y = (-9/2)x + 9 m = (-9/2) so a perpendicular slope is 2/9 y - 4 = (2/9)(x-1) y = (2/9)x - (2/9) + 4 y = (2/9)x - 38/9

OpenStudy (anonymous):

Ok thanks that makes sense... How do I change this into standard form then?